OpenStudy (anonymous):

If Sm is the sum of m consecutive natural numbers, what is the probability that Sm is a multiple of 21. (Given that m is less than or equal to 100.)

OpenStudy (anonymous):

Sm = m(m+1)/2, so to be divisible by 21, either m is a multiple of 21 or (m+1) is a multiply of 21

OpenStudy (anonymous):

idk @matricked

OpenStudy (anonymous):

@sourwing, the sequence does not have to start with 1.

OpenStudy (anonymous):

m = 21, 42, 63, 84, m + 1 = 21 m = 20 so, 20, 41, 62, 83,

OpenStudy (anonymous):

@eliassaab, since 0 isn't universal agreed that it is a natural number, I defined 0 to be not a natural number in this case.

OpenStudy (anonymous):

BUt the sequence can be 4, 5, 6

OpenStudy (anonymous):

P = 8/100

OpenStudy (anonymous):

Am I missing something?

OpenStudy (anonymous):

ah I see what you meant now. The sum starts at some number m. If this is the case, wouldn't this be not enough information?

OpenStudy (anonymous):

So I assume m starts at 1.

OpenStudy (anonymous):

But it does not say that in the problem. One cannot add assumptions about a question.

OpenStudy (anonymous):

well, true. Then I would say not enough information. We need know what m starts at, so we can eliminate the sum that starts at 1 and up to m.

OpenStudy (anonymous):

wrong questionah?

OpenStudy (anonymous):

not wrong, just not enough info. Unless you left something out.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

or m starts at 1 but the question did not explicitly stated so

OpenStudy (anonymous):

So I would say P = 8/100 = 4/25 anyway.

OpenStudy (anonymous):

idk whether it is correct or not

OpenStudy (anonymous):

Ten small boxes namely A, B, C.............J and each box can keep one ball only. Five big boxes X, Y, Z, W, U and each big box can keep ten balls maximum. A person is trying to transfer balls from small boxes to big boxes and at least 5 balls should be transferred. Find the number of ways of transfer.

OpenStudy (anonymous):

Please do not mix problems. Ask your problem in a new post. Returning the first problem. It has to specify the starting point. Suppose m=3, then we can get an infinite sequence of 3 consecutive numbers, like {1,23} {2,3,4} .... {100,101,102} {k,k+1,k+2}